• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ibrahim

    Log entry time 08:04:39 on June13,2009

    Entry number 280985

    keyword=Owl Shift Summary (June 13, 2009)


    Owl Shift Summary (June 13, 2009)
    ---------------------------------

    SL: Hassan Ibrahim
    TO: Miha Mihovilovic
    3rd: Lamiaa El Fassi

    00:00 shift started with cryo level problems on both left and right arms. Called the RC to check the problems and he is here now.

    01:25 The left arm is up now. The right dipole and Q3 are not coming back up because of some faults. The RC called the Tech on-call to check this.

    01:30 NMR 62.6%

    02:21 Restarted the HAPPEX DAQ. New run 31523.

    02:50 The Tech on-call is here. Going to restricted access.

    03:30 Fixed the problem. Going to beam permit while ramping up the right arm dipole.

    03:53 Restarted the HAPPEX DAQ. New run 31524. Started the Compton DAQ. New Run 17414.

    06:13 NMR 63.3%

    08:00 Shift ends with both arms look ok.



    Runs(LHRS Only):

    3529 C12 418k
    3530 He3 726k
    3531 He3 4.0M
    3532 He3 2.5M
    3533 C12 210k
    3534 He3 269k
    3535 He3 3.3M
    3536 He3 4.0M
    3537 He3 4.0M
    3538 He3 4.0M
    3539 He3 4.0M
    3540 He3 4.0M

    Runs (Both LHRS/RHRS)

    3541/22396 He3 4M/1M
    3542/22397 He3 still running at the end of the shift.